# DiffuGR: Generative Document Retrieval with Diffusion Language Models

### v4

@@ 5.1. Performance Comparison (Q1)
DDRO (PQ) 32.92 64.36 73.02 45.76
DDRO (TU) 38.24 66.46 74.01 50.07
Ours
LLaDA-Learnable 42.97 43.58 43.60 43.24
LLaDA-Linguistic 45.38 47.78 47.86 46.39
Dream-Learnable 43.60 44.22 44.29 43.88
Dream-Linguistic 45.05 47.05 47.21 45.96
Results on MS MARCO. To further validate the scalability and robustness of the proposed DiffuGR, we conducted experiments on the more complex MS MARCO dataset. Table 3 presents the performance comparison on the MS MARCO dataset, where DiffuGR still achieves impressive performance on R@1. However, DiffuGR’s performance on R@5 and R@10 is lower than that of some generative retrieval baselines. As for the issue already mentioned in the previous sections, this is related to the fact that the current diffusion language models do not support beam search-based top-$k$ generation. In addition, we also observe that most methods perform worse on the MS MARCO dataset than on NQ320K. This may be because the MS MARCO dataset is composed of real user search logs, while the documents in NQ320K are from Wikipedia. Therefore, MS MARCO is more aligned with real-world search scenarios and poses greater challenges.
5.2. Quality-Speed Tradeoff (Q2)
To answer Q2, we conduct an ablation study to evaluate the speed and quality with varying numbers of denoising iterations. Figure 3 shows the result on NQ320K. Results on MS MARCO show similar trends and can be found in Appendix A.2.1. As shown in Figure 3, increasing the number of iterations consistently improves retrieval performance, with higher R@1 scores achieved at larger iteration steps. However, the computational cost grows with the increase of iterations, leading to longer inference latency. For example, reducing the steps by half results in a noticeable decrease in latency while maintaining competitive retrieval accuracy, whereas further reducing the steps yields faster decoding at the expense of quality degradation. For comparison, we tested Qwen-2.5-7B R@1 and the number of samples processed per second. It can be observed that although Qwen achieved relatively good performance, its processing speed was relatively slow and failing to flexibly balance the tradeoff between performance and speed. These findings highlight a key advantage of DiffuGR, i.e., controllable generation process allows for dynamic adjustment of inference cost, making them suitable for scenarios with heterogeneous efficiency requirements. In particular, DiffuGR can leverage this flexibility to adaptively select iteration budgets depending on latency constraints of deployment.
@@ 5.2. Quality-Speed Tradeoff (Q2)
5.3. Effect of Pseudo Beam Search (Q3)
To answer Q3 and assess the effectiveness of the proposed pseudo beam search strategies, we conducted ablation experiments comparing three configurations of DiffuGR: (1) the vanilla model without pseudo beam search, (2) the model equipped with query augmentation, and (3) the model utilizing intermediate denoising candidates. All methods share identical architectures and training settings to ensure fair comparison.
@@ 5.3. Effect of Pseudo Beam Search (Q3)
Table 4. Effect of pseudo beam search on NQ320K. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. R@$k$ is short for Recall@$k$.
@@
LLaDA-Linguistic 66.15 66.90 66.90
+ query augmentation 66.62 66.96 66.96
+ intermediate denoising states 66.57 67.34 67.34
Dream-Linguistic 69.47 69.78 69.78
+ query augmentation 69.50 69.82 69.82
+ intermediate denoising states 69.57 69.83 69.83
Table 4 reports the Recall@{1,10,100} on NQ320K. Results on MS MARCO can be found in Appendix A.2.2. The results demonstrate that both strategies improve retrieval accuracy over the vanilla baseline. Specifically, query augmentation yields consistent gains by introducing diversity in query representations, allowing the model to explore multiple latent retrieval hypotheses.
The intermediate denoising approach also improves recall metrics, albeit to a lesser extent, as it relies on stochastic variations within the denoising process rather than semantic diversity. Although both methods have improved the performance of DiffuGR, the performance improvement is limited compared with the beam search technique under the auto-regressive model. Therefore, developing more effective methods to enable diffusion-based top-$k$ generation is an important direction for further improvement.
Table 5. Impact of denoising strategies on NQ320K. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. $\ddagger$ denotes the default setting. R@$k$ is short for Recall@$k$.
Method R@1 R@10 R@100
Dream-Linguistic
@@
5.4. Impact of Hyperparameters (Q4)
To answer Q4, experiments are conducted to investigate the effect of denoising strategies and model scales.
5.4.1. Impact of Denoising Strategies
Table 5 shows the impact of denoising strategies on NQ320K. Results on MS MARCO can be found in Appendix A.2.3. As shown in Table 5, different denoising strategies yield distinct retrieval performance under the same testing configuration. We observe that the maskgit plus and topk margin strategies consistently outperform the random baseline, demonstrating the benefit of confidence-guided denoising in reducing uncertainty accumulation during generation. The entropy strategy achieves comparable performance, suggesting that both local certainty (margin) and global uncertainty (entropy) are effective cues for progressive denoising. These findings indicate that token generation scheduling plays a crucial role in improving retrieval accuracy and stability of DiffuGR.
5.4.2. Impact of Model Scales
To investigate how model scales influence the performance of DiffuGR, we conducted a series of experiments using diffusion language models with varying parameter scales. Specifically, we experiment with Dream of different model depths, comprising 4, 8, and 16 transformer layers, derived from the original 28-layer Dream architecture. All models are trained until no further performance improvement is observed. To ensure fair comparisons, we keep all hyperparameters identical across models, including optimizer settings, learning rate schedule, and GPU configuration. Moreover, the smaller models are initialized from the corresponding layers of the pretrained Dream model; that is, the parameters of the first n layers are directly copied from the full 28-layer checkpoint. This strategy provides the smaller models with a strong initialization, enabling them to achieve better performance. Figure 4 presents the results on NQ320K. We can observe that larger models consistently outperform smaller ones on Recall@1, demonstrating that increasing model scales enhances the diffusion-based retrieval process. However, we also observe a significant increase in training time as model size grows, which means the training consumes more costs. In particular, the 16-layer model requires nearly twice the training resources of the 8-layer counterpart, reflecting a clear tradeoff between performance and efficiency. Overall, these findings indicates that moderately scaled models may strike a better balance between effectiveness and efficiency, making them more practical under limited resources.
Table 6. Perform comparison between diffusion-based generation and auto-regressive generation on the NQ320K dataset. R@$k$ is short for Recall@$k$. Best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level.
@@ 5.4.2. Impact of Model Scales
Qwen2.5-7B-Linguistic 68.96 69.27 69.27
Dream-Learnable 65.63 68.54 68.54
Dream-Linguistic 69.47 69.78 69.78
Besides, we also conducted experiments to assess the effectiveness of diffusion language models compared to state-of-the-art auto-regressive large language models under the same model size. Specifically, we benchmark our diffusion-based model against Qwen2.5-7B, a representative 7B-parameter auto-regressive model known for its strong performance across a variety of language understanding and generation tasks. This setup ensures that performance differences can be attributed to the generation paradigms rather than model scales. As shown in Table 6, our diffusion model achieves better retrieval accuracy compared to that of Qwen2.5-7B. Such results further verify the effectiveness of diffusion-based generation.
6. Conclusion
In this paper, we have proposed a generative document retrieval framework based on diffusion language models, dubbed DiffuGR, and have investigated its performance on two widely used datasets. Our results show that DiffuGR achieves competitive performance compared with strong auto-regressive baselines, while its generation paradigm is more suitable for generative document retrieval tasks. Moreover, the generation quality and speed of DocIDs can be flexibly controlled by adjusting the number of denoising iterations. However, DiffuGR still face several challenges: (1) DiffuGR does not support beam search, resulting in only marginal improvements in larger top-$k$ generation; (2) DiffuGR does not support constrained decoding, which may lead to invalidly generated DocIDs.
In future work, we plan to further improve the performance of diffusion language models for generative retrieval, aiming to address the above existing challenges. Besides, we are also interested in investigating whether larger diffusion models trained on more query-document pairs can further improve retrieval performance.
Appendix A Appendix
A.1. Baselines
The sparse retrieval baselines are as follows:
  • BM25 uses the tf-idf feature to measure term weights; we use the implementation from http://pyserini.io/.
  • DocT5Query expands a document with possible queries predicted by a finetuned T5 with this document as the input.
The dense retrieval baselines are as follows:
  • DPR (Karpukhin2020DensePR), a dual-encoder model using the representation of the [CLS] token of BERT.
  • ANCE (Xiong2020ApproximateNN), an asynchronously updated ANN indexer is utilized to mine hard negatives for training a RoBERTa-based dual-encoder model.
  • Sentence-T5 (Ni2021SentenceT5SS), a dual-encoder model that uses T5 to produce continuous sentence embeddings.
  • RepBERT (zhan2020repbert), a BERT-based model that generates fixed-length contextualized embeddings with query-document relevance computed via inner product similarity.
The generative retrieval baselines are as follows:
  • DSI (Tay2022TransformerMA), which represents documents using hierarchical $K$-means clustering results, and indexes documents using the first 32 tokens as pseudo-queries. As the original code is not open source, we refer to the results reported by (Wang2022ANC).
  • SEAL (Bevilacqua2022AutoregressiveSE) uses arbitrary n-grams in documents as DocIDs, and retrieves documents under the constraint of a pre-built FM-indexer. We refer to the results reported by (Wang2022ANC).
  • DSI-QG (Zhuang2022BridgingTG) uses a query generation model to augment the document collection. We refer to the results reported by (Wang2022ANC).
  • NCI (Wang2022ANC) uses a prefix-aware weight-adaptive decoder and various query generation strategies, including DocAsQuery and DocT5Query. In particular, NCI augments training data by generating 15 queries for each document.
  • Ultron (Zhou2022UltronAU) uses a three-stage training pipeline and represents documents through three types of identifiers.
  • ROGER (ROGER),which transfers knowledge from a dense retriever to a generative retriever via knowledge distillation.
  • MINDER (li2023multiview),which assigns multiple identifiers, including titles, n-grams, and synthetic queries, to documents and pairs them for indexing.
  • LTRGR (li2024learning),which trains on pairwise relevance objectives using margin-based ranking loss for optimization.
  • GenRRL (zhou2023enhancing),which incorporates pointwise, pairwise, and listwise relevance optimization through reinforcement learning, using document summaries and URLs as docids.
  • DDRO (mekonnen2025lightweight), which enhances generative retrieval by directly aligning generation with document-level relevance estimation.
A.2. Additional Results
Figure 5. Quality-speed tradeoff on MS MARCO.
A.2.1. Quality-Speed Tradeoff on MS MARCO
The results of quality-speed tradeoff on MS MARCO is shown in Figure 5.
Table 7. Effect of pseudo beam search on MS MARCO. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. R@$k$ is short for Recall@$k$.
Method R@1 R@10 R@100
LLaDA-Learnable 42.97 43.60 43.60
+ query augmentation 43.21 43.68 43.71
+ intermediate denoising states 43.59 43.89 43.89
Dream-Learnable 43.60 44.29 44.29
+ query augmentation 44.22 44.45 44.60
+ intermediate denoising states 44.20 44.50 44.61
A.2.2. Effect of Pseudo Beam Search on MS MARCO
The results of effect of pseudo beam search on MS MARCO is shown in Table 7.
Table 8. Impact of denoising strategies on MS MARCO. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. $\ddagger$ denotes the default setting. R@$k$ is short for Recall@$k$.
Method R@1 R@10 R@100
Dream-Linguistic
@@ A.2.2. Effect of Pseudo Beam Search on MS MARCO
A.2.3. Impact of Denoising Strategies on MS MARCO
The results of the impact of denoising strategies on MS MARCO is shown in Table 8

### v5

@@ 4.3. Implementation Details
DDRO (PQ)* 31.06 62.87 71.91 44.31
DDRO (TU)* 37.76 65.83 73.95 49.32
LLaDA-Learnable 43.42 63.79 65.20 52.29
LLaDA-Text 45.72 66.60 71.35 53.38
Dream-Learnable 46.83 63.48 64.49 55.15
Dream-Text 45.38 69.58 74.55 55.60
5. Results and Analysis
In this section, we present the experimental results to systematically evaluate the performance of DiffuGR and analyze the contribution of its core components. We list the following research questions to guide our experiments:
  1. RQ1
    How does DiffuGR perform compared to baseline methods?
  2. RQ2
    How does DiffuGR balance the quality–latency tradeoff?
  3. RQ3
    How does DiffuGR generate top-$k$ results with pseudo beam search?
  4. RQ4
5.1. Performance Comparison (RQ1)
Results on NQ320K and MS MARCO. Table 1 and Table 2 summarize the performances comparison on the NQ320K and MS MARCO, respectively. We draw the following observations: (1) The proposed DiffuGR significantly outperforms all generative retrieval baselines on two datasets. Specially, DiffuGR achieving 21.63 and 20.4 higher R@1 and MRR@10, respectively, compared to the best-performing baseline DDRO (Mekonnen et al., 2025) on NQ320K dataset. Even on the challenging MS MARCO dataset, DiffuGR also achieves favorable performance. This highlights the effectiveness of discrete diffusion paradigm in generative document retrieval task. (2) Different types of DocID affect the performance of DiffuGR. Overall, using text DocIDs yields better performance than using learnable DocIDs. This is likely because text-based DocIDs can directly leverage the semantic space of existing models, whereas learnable DocIDs require vocabulary expansion and semantic remapping, which further processes learning difficulty. Furthermore, we notice that different types of DocIDs exert a considerable impact on the performance of DiffuGR on the NQ320K dataset. We hypothesize that this may be attributed to the fact that the documents in the NQ320K dataset are equipped with well-formulated titles, which generally serve as a highly concise summary of the documents; thus, text-based DocIDs can yield better performance. Note that the focus of this paper lies in non-autoregressive generation of DocIDs other than the construction. (3) In addition, we also observe that most methods perform worse on the MS MARCO dataset than on NQ320K. This may be attributed to the fact that the MS MARCO dataset is a real web search dataset. For some reason, the titles of some documents may not summarize the document content well, and some documents even do not have a title at all. Furthermore, the queries provided by users also exhibit a certain degree of ambiguity and diversity, which impairs the performance.
Table 3. Zero-shot evaluation for BEIR dataset. The best results of generative retrieval are shown in bold. the second-best values of generative retrieval are underlined. Abbreviations denote DocIDs used for GR baselines: Naive - Naive String; Atomic - Atomic ID; SI - Semantic ID; PQ - Product Quantization; TU - Title + URL. * denotes that the results are reproduced using public code.
BEIR (nDCG@10)
Method Average Arg SciDocs
Sparse retrieval
BM25 21.95 29.1 14.8
DocT5Query 25.55 34.9 16.2
Generative retrieval
GENRE 0.30 0.0 0.6
DSI (Naive)* 0.10 0.1 0.1
DSI (Atomic)* 0.15 0.2 0.1
DSI (SI)* 3.85 1.8 5.9
NCI (SI)* 1.05 0.9 1.2
DDRO (PQ)* 0.03 0.04 0.02
DDRO (TU)* 0.55 1.0 0.1
LLaDA+Learnable 10.35 13.8 6.9
LLaDA+Text 9.25 13.1 5.4
Dream+Learnable 7.10 10.4 3.8
Dream+Text 4.10 6.2 2.0
Zero-shot performance on BEIR. Table 3 shows the zero-shot retrieval performance on the BEIR (Thakur et al., 2021) dataset after training on the NQ320K dataset. On the BEIR benchmark, DiffuGR consistently achieves competitive performance compared to other generative retrieval methods and attains the best average results. Its iterative denoising process generates high-confidence tokens step by step, which effectively reduces error accumulation and leads to better zero-shot generalization. In contrast, auto-regressive generative retrieval methods are more susceptible to cascading errors, resulting in inferior zero-shot performance. We further observe that generative retrieval methods based on semantic IDs exhibit relatively better zero-shot performance. We hypothesize that this may be due to the short length of semantic IDs, which reduces the retrieval difficulty of new documents. However, sparse retrieval methods generally outperform generative retrieval models in the zero-shot setting, highlighting generalization as a key challenge of generative retrieval.
Table 4. Evaluation of catastrophic forgetting without beam search. The best results are shown in bold. The second-best values are underlined. Abbreviations denote DocIDs used for GR baselines: SI – Semantic ID; PQ – Product Quanti- zation; TU – Title + URL. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. * denotes that the results are reproduced using public code.
$\text{NQ320K}_{old}$
Method R@1 R@10 MRR@10 nDCG@10
Training on $\text{NQ320K}_{old}$
DSI (SI)* 29.31 33.58 32.96 32.87
DDRO (TU)* 43.82 44.35 44.17 44.21
DDRO (PQ)* 49.97 51.53 51.32 51.28
LLaDA+Learnable 66.02 69.01 67.44 67.85
LLaDA+Text 69.32 69.59 69.43 69.47
Dream+Learnable 65.80 68.76 67.20 67.60
Dream+Text 69.57 69.77 69.65 69.65
Continual Training on $\text{NQ320K}_{new}$
DSI (SI)* 11.53 13.98 13.83 13.23
DDRO (TU)* 14.98 15.72 15.64 15.71
DDRO (PQ)* 16.45 17.91 17.57 17.60
LLaDA+Learnable 19.95 20.87 20.38 20.51
LLaDA+Text 39.72 39.87 39.78 39.81
Dream+Learnable 22.06 23.15 22.58 22.73
Dream+Text 45.12 45.27 45.18 45.21
Catastrophic forgetting problem. To investigate whether catastrophic forgetting occurs when applying continual learning to adapt DiffuGR to new documents, we randomly select 10000 documents from NQ320K to construct a new dataset $\text{NQ320K}_{new}$. The rest of the documents are then labeled $\text{NQ320K}_{old}$. As shown in Table 4, after continual training on the $\text{NQ320K}_{new}$, the performance of generative retrieval methods on the $\text{NQ320K}_{old}$ significantly declined. It suggests that these methods encounter a distinct catastrophic forgetting issue. This suggests that merely conducting continuous training on new documents for generative retrieval does not represent an optimal approach. This is a common challenge in the field of generative retrieval. The approach to tackle this challenge lies outside the purview of this paper. We thus designate it as future work. However, DiffuGR still maintains competitive performance. We hypothesize that bidirectional attention and non-autoregressive generation mitigate the impact of catastrophic forgetting on DiffuGR. Furthermore, we observe that the DiffuGR with text-based DocID demonstrates superior anti-catastrophic forgetting performance compared with the learnable DocID counterpart, as evidenced by the reduced magnitude of performance decline of DiffuGR following continuous learning. We hypothesize that this may be attributed to the higher lexical diversity of text-based DocIDs, which reduces the risk of being overwritten during continuous learning.
@@ 5.1. Performance Comparison (RQ1)
Figure 4. Quality-speed tradeoff on MS MARCO. By adjusting the denoising steps, the performance of DiffuGR can be flexibly tuned towards either speed or quality.
5.2. Quality-Speed Tradeoff (RQ2)
To answer RQ2, we conducted an ablation study to evaluate the speed and quality with varying numbers of denoising iterations. Figure 3 and Figure  4 show the results on NQ320K and MS MARCO, respectively. As shown in Figure 3 and Figure 4, increasing the number of iterations consistently improves DiffuGR performance, with higher R@1 scores achieved at larger iteration steps. However, the computational cost grows with the increase of iterations, leading to longer inference latency. For example, reducing the steps by half results in a noticeable decrease in latency while maintaining competitive retrieval accuracy, whereas further reducing the steps yields faster decoding at the expense of quality degradation. For comparison, we tested Qwen-2.5-7B R@1 and the number of samples processed per second on NQ320k. It can be observed that although Qwen achieved relatively competitive performance, its speed is relatively slow and it fails to flexibly balance the tradeoff between performance and speed. These findings underscore a crucial advantage of DiffuGR: its controllable generation process enables the dynamic adjustment of inference cost, rendering it appropriate for scenarios characterized by efficiency requirements.
@@ 5.2. Quality-Speed Tradeoff (RQ2)
Table 5. Effect of pseudo beam search on NQ320K. CTE is short for candidate trajectory extraction. BS is short for beam search. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. R@$k$ is short for Recall@$k$.
Method R@1 R@10 R@100 MRR@10
LLaDA-Text 66.15 66.90 66.90 66.26
+ query aug 66.62 66.96 66.96 66.33
+ CTE 66.57 67.34 67.34 66.97
+ diffusion BS 66.29 81.25 84.45 71.86
Dream-Text 69.47 69.78 69.78 69.57
+ query aug 69.50 69.82 69.82 69.71
+ CTE 69.57 69.83 69.83 69.72
+ diffusion BS 69.73 82.82 85.83 74.65
Table 6. Effect of pseudo beam search on MS MARCO. CTE is short for candidate trajectory extraction. BS is short for beam search. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. R@$k$ is short for Recall@$k$.
Method R@1 R@10 R@100 MRR@10
LLaDA-Learnable 42.97 43.60 43.60 43.24
+ query aug 43.21 43.68 43.71 43.26
+ CTE 43.59 43.89 43.89 43.27
+ diffusion BS 43.42 65.20 65.68 52.29
Dream-Learnable 46.60 46.92 46.92 45.96
+ query aug 46.89 46.95 46.98 45.97
+ CTE 46.62 46.93 46.92 45.97
+ diffusion BS 46.83 64.49 64.70 55.15
5.3. Effect of Pseudo Beam Search (RQ3)
To answer RQ3 and assess the effectiveness of the proposed pseudo beam search strategies, experiments are conducted on DiffuGR with the three beam search strategies. Table 5 and Table 6 report the Recall@{1,10,100} and MRR@10 on NQ320K and MS MARCO, respectively. The results indicate that three strategies enhance retrieval accuracy compared to the vanilla baseline. However, distinct methods lead to significantly different degrees of performance improvement: (1) Query augmentation method yields marginal gains by introducing diversity in query representations. We find the reason lies in the fact that most of the DocIDs generated by this method are duplicates. Duplicate DocIDs result from the trait of high-performance generative retrieval models: they produce identical DocIDs for semantically identical queries with different expressions or wordings. We hypothesize that this method could effectively enhance model performance when dealing with ambiguous queries. This is because query augmentation can mitigate the ambiguity in queries, thereby improving the model’s top‑k performance. However, research on ambiguous queries falls beyond the scope of this work, and we defer it to future studies. (2) The Candidate Trajectory Extraction (CTE) method enhances the top‑k performance of DiffuGR by leveraging intermediate results from the denoising process. However, its improvements on the R@10/100 and MRR@10 metrics remain limited. This limitation stems from two main factors. First, DocIDs produced in denoising are noisy. Second, Once a correct DocID is generated early in decoding, the model tends to retain it in later steps, causing substantial repetition of DocIDs. A key advantage of the method is that it introduces no additional computational overhead during inference. Since it only requires retaining the outputs from the forward denoising process, it remains applicable even under resource‑constrained inference scenarios. (3) The Diffusion Beam Search method achieves significant performance improvement on the R@10/100 and MRR@10 metrics. It is important to note, however, that this method requires substantial computational resources, primarily for two reasons. First, it expands the number of inference samples by a factor of B, thereby increasing computational costs by approximately B times. Second, the selection process for each token involves applying a top-k algorithm, which introduces additional computational overhead.
Table 7. Impact of denoising strategies on NQ320K without beam search. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. $\ddagger$ denotes the default setting. R@$k$ is short for Recall@$k$.
@@ 5.3. Effect of Pseudo Beam Search (RQ3)
Dream-Text
@@
Table 8. Impact of denoising strategies on MS MARCO without beam search. The best results are shown in bold. $\dagger$ indicates the result is significantly improved with paired $t$-test at $p<0.05$ level. $\ddagger$ denotes the default setting. R@$k$ is short for Recall@$k$.
@@
Dream-Text
@@
5.4. Impact of Hyperparameters (RQ4)
5.4.1. Impact of Denoising Strategies
Table 7 and Table 8 show the impact of denoising strategies on NQ320K and MS MARCO, respectively. As shown in Table 7 and Table 8, different denoising strategies yield distinct retrieval performance under the same testing configuration. We observe that the maskgit plus and topk margin strategies consistently outperform the random baseline, demonstrating the benefit of confidence-guided denoising in reducing uncertainty accumulation during generation. The entropy strategy achieves comparable performance, suggesting that both local certainty (margin) and global uncertainty (entropy) are effective cues for progressive denoising. In practical applications, these strategies can be adopted to provide users with two types of query styles: precision-oriented and diversity-oriented.
5.4.2. Impact of Model Scales
In order to explore the impact of model scales on the performance of DiffuGR, a series of experiments were carried out, in which diffusion language models with different parameter scales were utilized. Specifically, we experiment with Dream of different model depths, comprising 4, 8, and 16 transformer layers, derived from the original 28-layer Dream architecture. Figure 5 presents the results on NQ320K. We can observe that larger models consistently outperform smaller ones on R@1, demonstrating that increasing model scales enhances the diffusion-based retrieval process. However, we also observe a significant increase in training time as model size grows, which means training consumes more costs.
Table 9. Perform comparison between diffusion-based generation and autoregressive generation on the NQ320K dataset without beam search. R@$k$ is short for Recall@$k$. Best results are shown in bold.
Method R@1 R@10 R@100
Qwen2.5-7B-Text 68.96 69.27 69.27
Dream-Learnable 65.63 68.54 68.54
Dream-Text 69.73 69.78 69.78
Besides, we also conducted experiments to assess the effectiveness of diffusion language models compared to auto-regressive large language models under the same model size. Specifically, we benchmark our diffusion-based model against Qwen2.5-7B. This setup ensures that performance differences can be attributed to the generation paradigms rather than model scales. As shown in Table 9, Dream+Text achieves better R@1 compared to that of Qwen2.5-7B. The results indicate that, within the same parameter scale, DiffuGR is comparable with auto-regressive models. Moreover, it offers a novel research direction for generative retrieval.
6. Conclusion
In this paper, we have proposed a generative document retrieval framework based on diffusion language models, termed DiffuGR. DiffuGR is compatible with two different types of DocIDs: learnable DocID and text-based DocID. DiffuGR employs masked diffusion language models to generate DocIDs for a given query. It provides a forward process that progres- sively masks DocID tokens and learns a reverse denoising process that iteratively reconstructs the complete identifier sequence. We have conducted extensive experiments to examine the performance of DiffuGR on two widely used benchmarks. Our results show that DiffuGR achieves outstanding performance compared with strong auto-regressive baselines, while its generation paradigm is more suitable for generative document retrieval tasks. Moreover, the generation quality and latency of DocIDs can be flexibly controlled by adjusting the number of denoising iterations.
The limitations of this work are threefold. First, we do not optimize DocID construction, as our primary focus is on non-autoregressive DocID generation. Second, DiffuGR currently does not support constrained decoding, which may result in the generation of invalid DocIDs. Third, DiffuGR still suffers from catastrophic forgetting under continual training, which is also shared by existing generative retrieval methods. Addressing these limitations constitutes important directions for future work. In particular, we plan to explore diffusion-aware DocID construction strategies and validity-preserving decoding mechanisms, as well as improve the robustness of diffusion-based generative retrievers under dynamic and continually evolving corpora. We hope this work will provide useful insights and serve as a foundation for future research in generative retrieval.